Use the information below to answer the questions that follow.
U.S. $ EQUIVALENT CURRENCY PER U.S. $
  U.K. pound (£) 1.5542             .6434            
  Canada dollar (Can$) 1.0251           .9755              
Which would you rather have, $100 or £100?
(Click to select)100 Dollars100 Pounds
Which would you rather have, $100 Canadian or £100?
(Click to select)100 Canadian Dollars100 Pounds

What is the cross-rate for Canadian dollars in terms of British pounds? (Do not include the Canadian dollar sign, Can$. Round your answer to 4 decimal places, e.g., 32.1616.)

  Cross-rate Can$  /£

What is the cross-rate for British pounds in terms of Canadian dollars? (Do not include the pound sign, £. Round your answer to 4 decimal places, e.g., 32.1616.)

  Cross-rate £  /Can$
Answer #1

The answer is Pound 100 since value of Pound is higher than Dollar

The answer is Pound 100 since value of Pound is higher

Cross rate Can$/pound = (can $/dollar)*(dollar/pound)

= (0.9755)*1.5542

= can$1.516122/pound

Pound/can dollar = 0.6434*1.0251 = 0.659549 pound/can dollar